WAC Notes July 18 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << Instructions

WAC Notes July 17 2019 << >> WAC Notes July 19 2019

July 18 2019 all 3 Shifts

Slug 5: 70 uA production - Test Swing Shift

Data was recovered from 3424, 3429, and 3430 current ramp runs with better beam current lower limit cut

  • Bad/Current Ramp: 3424 - Aq FB was insufficient as this was a current ramp
  • Good: 3425
  • Good: 3426
  • Good: 3427
  • Good: 3428
  • Bad/Current Ramp: 3429 - Aq FB was insufficient as this was a current ramp
  • Bad/Current Ramp: 3430 - Aq FB was insufficient as this was a current ramp
  • Calibration: 3431
  • Calibration: 3432

Slug 6: 50 uA Production - Stable Owl Shift

Lots of production, only 1 major problem (lost run 3533 with 9k muls due to blinder fail from EPICs fastlogger and Compton IOC)

There is a gap here in run number, as the intervening PQB studies flipped the HWP state without being accounted properly the first time (and the following IHWP out runs are 1 arm only)

  • Good: 3450 - Good after BPM cuts
  • Good: 3451 - Good after BPM cuts
  • Good: 3452 - Good after BPM cuts

Slug 7: 50 uA Production - During PQB Beam Studies

  • Good: 3444
  • Good: 3445
  • Good: 3446 - double difference in BCM returns with vengence


Slug 8: 50 uA Production - Truncated Swing Shift

Caryn changed the IHWP state to IN

  • Good: 3453 - RQ2 tripped off
  • Good: 3454
  • Good: 3455
  • Next: access, and PQB to follow at low current, then Moller, then DAQ, then another access for magnets
  • Runs 3456 -> are slug 8

End of slug 7 -> Removed S0 -> should have a tiny ~1% effect on widths now

Non-Production Runs and Bad Runs

Pedestal/Calibrations Updates

In order to salvage current ramps and trips on large chunks of data and save time/save the regression/pedestal usefulness:

  • Change the lower limit current cut to be nominal-15 uA (35 for 50 uA, 55 for 70 uA)
  • Updating the event cuts to have nominal-15 uA beam current lower cut Starting at run 3456 (now) and also setting specifically for 3 70 uA current ramp runs 3424, 3429, 3430 (suspect Aq, but worth a try)
    • 50 uA: prexCH_beamline_eventcuts.2948-.map -> prexCH_beamline_eventcuts.3456-.map (pending beam...)
      • 3451: BCM cut fixed, but PQB efforts messed with BPMs too much
      • 3452: BCM cut fixed, but PQB efforts messed with BPMs too much
      • 3453: Fixed
    • 70 uA: prexCH_beamline_eventcuts.2948-.map -> prexCH_beamline_eventcuts.[3424,3429,3430].map
    • Follow up - cut on current at nominal - 15 uA -> applied to current ramp runs
      • 3424: Cuts out half the prior data... marginal improvements
      • 3429: No effect on anything (since the beam was always at 60-70 anyway)
      • 3430: Cuts out 2/3rds of the prior passing data - fixes regression!
  • Caryn updated the Main Det and SAM pedestals from the current scans, for 50 uA and 70 uA data:
    • Here, noted Here
    • New Pedestal Map for 70 uA 3424-3430 => prex_sam_pedestal.3424-3430.map, prex_maindet_pedestal.3424-3430.map
    • New Pedestal Map for 50 uA 3456- => prex_sam_pedestal.3456-.map, prex_maindet_pedestal.3456-.map
    • Also changed retroactively from 3398- the prex_sam_pedestal.3398-.map to match prex_sam_pedestal.3456-.map 50 uA pedestals

Issues

  • Needed nominal - 15 uA BCM lower limit cut.
  • Need BPM stability/burp/holdoff kind of cut to save some strange noise in 3403, 3404, and some PQB runs this afternoon.
  • It looks like there is non-symmetry between arms (before RQ2 failed)

Other Status Updates

  • PQB runs in the afternoon need more in depth (BPM cut) analyzing - see notes
  • S0 Removed - midway through slug 8... but at least we know when (run 3456 is the production run it happens after)